OpenStudy (anonymous):

how do you divide 42 2/9 by 6 6/7?

OpenStudy (anonymous):

it will be: 380/9 devided by 48/7 just take the reciprocal of the denominator then multiply: 380/9 x 7/48 = 2660/432 = 323/54 or 6 1/54

OpenStudy (anonymous):

\[42+2/9 = (42 \times9+2)/9 = (378+2)/9 = 380/9\] \[6+6/7=(6 \times7 +6)/7=48/7\]

OpenStudy (anonymous):

The correct answer is:\[\frac{665}{108} \]\[\frac{2660}{432}=\frac{665*4}{108*4} \]Greatest Common Divisor of 665 and 108 is 1.